OpenStudy (anonymous):

Solve the equation., and provide x values where 0 < x < 3pi: cos^2 x + cos(2x) = 5/4

myininaya (myininaya):

\[\cos(2x)=\cos^2(x)-\sin^2(x)=\cos^2(x)-(1-\cos^2(x))=2\cos^2(x)-1\] Use this identity

OpenStudy (campbell_st):

Use the substitution cos(2x) = 2cos^2(x) - 1 then cos^2(x) + 2cos^2(x) - 1 = 5/4 or 3cos^2(x) = 9/4 cos^2(x) = 3/4 \[\cos(x) = \pm \sqrt{3}/2\] \[\cos^{-1}(\sqrt{3}/2) = \pi/3\] you'll need to find the rest of the values of x over the given domain